OpenStudy (anonymous):

solve the separable DE dx/dt -x^3=x

OpenStudy (anonymous):

do you want to solve for x or find the derivative?

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

well wait...its separable equations so we need to get dx/dt by itself and then take the integral of both sides. The right side calls for partial fractions which is what I am having trouble with, then we solve for x

OpenStudy (anonymous):

the farthest I got is \[\int\limits dt=\int\limits \frac{ 1 }{ x+x^3}\]

OpenStudy (anonymous):

\[\begin{align*}\frac{dx}{dt}-x^3&=x\\\\ \frac{dx}{dt}&=x^3+x\\\\ \frac{dx}{x^3+x}&=dt\\\\ \int\frac{dx}{x^3+x}&=\int dt\\\\ \int\frac{dx}{x(x^2+1)}&=\int dt \end{align*}\] Partial fractions, or trig sub. I prefer the former: \[\begin{align*}\frac{1}{x(x^2+1)}&=\frac{A}{x}+\frac{Bx+C}{x^2+1}\\\\ 1&=A(x^2+1)+x(Bx+C)\\ 1&=Ax^2+A+Bx^2+Cx\end{align*}~~\implies~~\begin{cases}A+B=0\\C=0\\A=1\end{cases}\] which gives \(A=1,~B=-1,~C=0\). \[\begin{align*}\int\frac{dx}{x(x^2+1)}&=\int dt\\\\ \int\left(\frac{1}{x}-\frac{1}{x^2+1}\right)~dx&=\int dt \end{align*}\]

OpenStudy (anonymous):

thank you! can you go into detail on how you found A, B and C?

OpenStudy (anonymous):

\[\begin{align*}\frac{1}{x(x^2+1)}&=\frac{A}{x}+\frac{Bx+C}{x^2+1}\\\\ \frac{1}{x(x^2+1)}&=\frac{A(x^2+1)}{x(x^2+1)}+\frac{x(Bx+C)}{x(x^2+1)}\\\\ \frac{1}{x(x^2+1)}&=\frac{A(x^2+1)}{x(x^2+1)}+\frac{x(Bx+C)}{x(x^2+1)}\\\\ 1&=A(x^2+1)+x(Bx+C)\\\\ 1&=Ax^2+A+Bx^2+Cx\\\\ 1&=(A+B)x^2+Cx+A\end{align*}\] In the last line, we group the like terms (the equivalent powers of \(x\)). From here, we match up the coefficients of the powers of \(x\) on both sides: \[0x^2+0x+1=(A+B)x^2+Cx+A\] which gives \[\begin{cases}A+B=0\\C=0\\A=1\end{cases}\]